# Compatibility of Drinfeld presentations for split affine Kac–Moody quantum symmetric pairs

### v1

@@ 9.1. $q$-characters of quantum affine algebras
Fix $\mathbf{c}=(c_{0},\cdots,c_{n})\in{\mathbb{C}}^{n+1}$ and let $C$ be the image of $\mathfrak{C}$ in $\mathbf{U}^{\imath}_{\mathbf{c}}$. Let $\operatorname{Rep}\mathbf{U}$ be the monoidal category of finite dimensional representations of $\mathbf{U}$. It acts on $\operatorname{Rep}\mathbf{U}^{\imath}_{\mathbf{c}}$, the category of finite dimensional representations of $\mathbf{U}^{\imath}_{\mathbf{c}}$, via the coproduct $\Delta_{\mathbf{c},\mathbf{s}}$ (2.29). It follows from [DK19, Sec. 3.5] (see also [Prz23, Lemmas 2.4]) that any such monoidal action is, up to twisting by a character, equal to the “standard” monoidal action with $\mathbf{s}=(0,\cdots,0)$. Therefore, henceforth we shall assume that we are working with the standard monoidal action.
We will use the same notations for the Grothendieck groups of the categories above. From this point of view, $\operatorname{Rep}\mathbf{U}$ is a ring acting on $\operatorname{Rep}\mathbf{U}^{\imath}_{\mathbf{c}}$. Let $U_{q}(\widetilde{\mathfrak{h}})$ be the subalgebra of $\mathbf{U}$ generated by $h_{i,k}$ ($i\in\mathbb{I}$, $k<0$). Following [FR99, FM01], set
(9.1) $Y_{i,a}=$ $\ K_{\omega_{i}}^{-1}\exp\left(\sum_{k>0}\tilde{h}_{i,-k}a^{k}z^{k}\right)\in{U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]\qquad(a\in{\mathbb{C}}^{\times}),$
@@
and $\widetilde{C}(q)$ is the inverse of the $q$-Cartan matrix. Let $\mathcal{Y}={\mathbb{Z}}[Y_{i,a}^{\pm 1}]_{i\in I,a\in{\mathbb{C}}^{\times}}$. By [FR99, Theorem 3], there exists an injective ring homomorphism $\chi_{q}\colon\operatorname{Rep}\mathbf{U}\to\mathcal{Y}\subset{U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$, called the $q$-character map, given by
@@
The expression above derives from the Khoroshkin-Tolstoy-Levendorsky-Soibelman-Stukopkin-Damiani (KTLSSD) factorization of the universal $R$-matrix [KT94, LSS93, Dam98]. However, by [FM01, Proposition 2.4], the $q$-character map can equivalently be defined more explicitly in terms of the joint spectrum of the Drinfeld-Cartan operators (i.e., the coefficients of the series $\boldsymbol{\phi}^{\pm}_{i}(z)$). More precisely, there is a one-to-one correspondence between the monomials occurring in $\chi_{q}(V)$ and the common eigenvalues of $\boldsymbol{\phi}^{\pm}_{i}(z)$ on $V$. Let $V=\bigoplus_{\gamma}V_{\gamma}$, with $\gamma=(\gamma^{\pm}_{i,\pm m})_{i\in I,m\in{\mathbb{Z}}_{\geq 0}}$, where
@@ 9.2. Boundary $q$-characters
A two-factor universal $K$-matrix for affine quantum symmetric pairs was defined in [AV24]. However, a KTLSSD-type factorization is not yet available. [^7] In the finite case, a factorization of quasi $K$-matrices was established in [DK19, WZ23]. Therefore, we propose to initiate the study of $q$-characters for affine quantum symmetric pairs based on the Lu–Wang (Drinfeld-type) presentation.
@@
$\mathcal{K}^{0}=\frac{1-q_{i}^{-2}Cz^{2}}{1-Cz^{2}}\exp\left(-(q-q^{-1})\sum_{i\in\mathbb{I}_{0}}\sum_{k>0}\frac{k}{[k]_{q_{i}}}H_{i,k}\otimes\tilde{h}_{i,-k}z^{k}\right)\in{\mathbf{U}^{\imath}_{\mathbf{c}}\otimes U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace].$

### v2

@@ 9.1. $q$-characters of quantum affine algebras
Fix $\mathbf{c}=(c_{0},\cdots,c_{n})\in({\mathbb{C}}^{\times})^{n+1}$ and let $C$ be the image of $\mathfrak{C}$ in $\mathbf{U}^{\imath}_{\mathbf{c}}$. Let $\operatorname{Rep}\mathbf{U}$ be the monoidal category of finite dimensional representations of $\mathbf{U}$. It acts on $\operatorname{Rep}\mathbf{U}^{\imath}_{\mathbf{c}}$, the category of finite dimensional representations of $\mathbf{U}^{\imath}_{\mathbf{c}}$, via the coproduct $\Delta_{\mathbf{c},\mathbf{s}}$ (2.30). It follows from [DK19, Sec. 3.5] (see also [Prz23, Lemmas 2.4]) that any such monoidal action is, up to twisting by a character, equal to the “standard” monoidal action with $\mathbf{s}=(0,\cdots,0)$. Therefore, henceforth we shall assume that we are working with the standard monoidal action.
We will use the same notations for the Grothendieck groups of the categories above. From this point of view, $\operatorname{Rep}\mathbf{U}$ is a ring acting on $\operatorname{Rep}\mathbf{U}^{\imath}_{\mathbf{c}}$. Let $U_{q}(\widetilde{\mathfrak{h}})$ be the subalgebra of $\mathbf{U}$ generated by $h_{i,k}$ ($i\in\mathbb{I}_{0}$, $k<0$). Following [FR99, FM01], set
(9.1) $Y_{i,a}=$ $\ K_{\omega_{i}}^{-1}\exp\left(-(q-q^{-1})\sum_{k>0}\tilde{h}_{i,-k}a^{k}z^{k}\right)\in{U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]\qquad(a\in{\mathbb{C}}^{\times}),$
@@
and $\widetilde{C}(q)$ is the inverse of the $q$-Cartan matrix. Let $\mathcal{Y}={\mathbb{Z}}[Y_{i,a}^{\pm 1}]_{i\in\mathbb{I}_{0},a\in{\mathbb{C}}^{\times}}$. By [FR99, Theorem 3], there exists an injective ring homomorphism $\chi_{q}\colon\operatorname{Rep}\mathbf{U}\to\mathcal{Y}\subset{U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$, called the $q$-character map, given by
@@
The expression above derives from the Khoroshkin-Tolstoy-Levendorsky-Soibelman-Stukopkin-Damiani (KTLSSD) factorization of the universal $R$-matrix [KT94, LSS93, Dam98]. However, by [FM01, Proposition 2.4], the $q$-character map can equivalently be defined more explicitly in terms of the joint spectrum of the Drinfeld-Cartan operators (i.e., the coefficients of the series $\boldsymbol{\phi}^{\pm}_{i}(z)$). More precisely, there is a one-to-one correspondence between the monomials occurring in $\chi_{q}(V)$ and the common eigenvalues of $\boldsymbol{\phi}^{\pm}_{i}(z)$ on $V$. Let $V=\bigoplus_{\gamma}V_{\gamma}$, with $\gamma=(\gamma^{\pm}_{i,\pm m})_{i\in\mathbb{I}_{0},m\in{\mathbb{Z}}_{\geq 0}}$, where
@@ 9.2. Boundary $q$-characters
The notion of a quasi-$K$-matrix appeared first in the work of Bao and Wang [BW18a, BW18b] on canonical bases for quantum symmetric pairs, as an intertwiner between the bar involutions on the quantum group and the coideal subalgebra. This construction of the quasi-$K$-matrix was later extended to arbitrary Kac–Moody type by Balagović and Kolb in [BK19] and, in finite type, led to the realization of the universal $K$-matrix as a coideal intertwiner [BW18b, BK19]. Later, Appel and Vlaar [AV22b, AV24] generalized these methods to obtain a universal $K$-matrix in the affine case. However, a KTLSSD-type factorization is not yet available. [^10] In the finite case, a factorization of quasi $K$-matrices was established in [DK19, WZ23]. Therefore, we propose to initiate the study of $q$-characters for affine quantum symmetric pairs based on the Lu–Wang (Drinfeld-type) presentation.
@@
$\mathcal{K}^{0}=\frac{1-q_{i}^{-2}\mathfrak{C}z^{2}}{1-\mathfrak{C}z^{2}}\exp\left(-(q-q^{-1})\sum_{i\in\mathbb{I}_{0}}\sum_{k>0}\frac{k}{[k]_{q_{i}}}H_{i,k}\otimes\tilde{h}_{i,-k}z^{k}\right)\in{\mathbf{U}^{\imath}_{\mathbf{c}}\otimes U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace].$
@@
Conflict of interest statement
The authors have no competing interests to declare that are relevant to the content of this article.
Data availability statement
The authors declare that all the data supporting the findings of this article are available within the paper.